OpenStudy (anonymous):

need help. solve for x. \[e ^{x ^{2}-1}=1/x\]

OpenStudy (anonymous):

Are you allowed to use Logarithms??

OpenStudy (anonymous):

Taking ln both the sides: \[\ln_e(e^{x^2-1}) = \ln_e(\frac{1}{x})\] \[\ln_e(e^{x^2-1}) = \ln_e(1) - \ln_e(x)\] \[(x^2 - 1)(1) = 0 - \ln_e(x)\] \[1 - x^2 = \ln_e(x)\] Got stuck now.

OpenStudy (anonymous):

we know that x>0 for x>1 \( e^{x^2-1}>1 \ \ and \ \ \frac{1}{x}<1 \) for x<1 \( e^{x^2-1}<1 \ \ and \ \ \frac{1}{x}>1 \) so x=1

OpenStudy (anonymous):

yeah i think my teacher would like us to use logs

OpenStudy (anonymous):

@waterineyes use the same approach for ur solution

OpenStudy (anonymous):

Yeah for x = 1, \(1- (1)^2 = ln_e(1)\) \(0 = 0\) Yeah x = 1 can be one possibility..

OpenStudy (anonymous):

And I think there is one possibility only... So, \[\large \color{green}{x = 1}\]

OpenStudy (anonymous):

ok thank you for your help!

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

Say thanks to mukushla also @jkd13

OpenStudy (anonymous):

it was a thanks to both of you :)

OpenStudy (anonymous):

the method used here is the method i'd use but there's a problem... you'll end up with like what waterineye got and that is an \(\large x^2 \) term and a \(\large lnx \) term. the problem here is to isolate the x and there is no elematary algebraic method that i know of to do that.....

OpenStudy (anonymous):

ok thank you cause thats whats throwing me off but i think this will be one to ask my instructor to see what he's looking for.

OpenStudy (anonymous):

yw...:)